OpenStudy (anonymous):

Tough probability question: If you split up the word ELEMENT and put the letters back in a random order, what is the probability that the three Es are consecutive?

OpenStudy (anonymous):

This was at MathCounts yesterday. I got an answer, but I'm not sure if it is right. I got 1/588

OpenStudy (anonymous):

there are probably a couple different ways to do this, and probably some easier than this, but you can figure out the probability that they are picked first, that would be \[\frac{3}{7}\times \frac{2}{6}\times \frac{1}{5}\]

OpenStudy (anonymous):

That's not what I'm asking

OpenStudy (anonymous):

\[\frac{1}{42}\]

OpenStudy (anonymous):

yes, but then we can look at the possible placement of the three E EEE X X X X X EEE X X X XX EEE X X XXX EEE X XXXX EEE

OpenStudy (jamesj):

Suppose the three Es are first, i.e., we have EEExxxx How many such possibilities are there? 3! 4! Now we can also have xEEExxx, another 3! 4! xxEEExx, 3!4! xxxEEEx, 3!4! xxxxEEE, 3!4! for a total of 5 . 3!4! Now divide that by the total number of possible arrangements.

OpenStudy (anonymous):

since each of these has the same probability, and since each of these is different (disjoint events) then you can add them up

OpenStudy (anonymous):

Wait a sec there, they are NOT the same EEExxxx has a different probability than xxEEExx

OpenStudy (jamesj):

But there are an equal number of possible arrangements EEExxxx and xxEEExx

OpenStudy (anonymous):

xxEEExx = x x 3/5 2/4 1/3 x x

OpenStudy (anonymous):

as i said, probably several ways to do this. i get \[5\times \frac{3}{7}\times \frac{2}{6}\times \frac{1}{5}\] \[=\frac{1}{7}\] but i could certainly by wrong

OpenStudy (jamesj):

Using my method I arrive at exactly the same answer: \[ \frac{ 5 \cdot 3! 4!}{7!} = \frac{1}{7} \]

OpenStudy (anonymous):

like is said, more ways that one to skin a cat

OpenStudy (anonymous):

do you still consider the arrangements of the e's?

OpenStudy (anonymous):

skin a cat? What I did was I knew that there was 5 possiblities of three consecutive e's

OpenStudy (anonymous):

EEExxxx has a different probability than xxEEExx \[\frac{3}{7}\times \frac{2}{6}\times \frac{1}{5}\times \frac{4}{4}\times \frac{3}{3}\times \frac{2}{2}\times \frac{1}{1}\] same as \[\frac{3}{7}\times \frac{2}{6}\times \frac{1}{5}\]

OpenStudy (anonymous):

also, the total number of possibilities of letter combinations is not 7!, it's 7!/3! because there are 3 Es

OpenStudy (jamesj):

In my calculation, I am considering the three Es to be distinguishable, in which case there are indeed 7! possibilities. That is why is each of the scenarios EEExxxx, etc. I also count the 3! = 6 possible arrangements of the three Es.

OpenStudy (anonymous):

but EEExxxx is not different from EEExxxx

OpenStudy (jamesj):

Right, but I am counting correctly. That's all.

OpenStudy (anonymous):

now we try xxEEExx \[\frac{4}{7}\times \frac{3}{6}\times \frac{3}{5}\times \frac{2}{4}\times \frac{1}{3}\times \frac{2}{2}\times \frac{1}{1}\] think it is identical

OpenStudy (anonymous):

I don't see where you're going james

OpenStudy (jamesj):

In other words, \[ E_1E_2E_3xxxx \] is not the same as \[ E_3E_1E_2xxxx \] But it is a possible arrangement. So I am counting up all of the possible arrangements of the seven letters which give three Es in a row. That is \[ 5 \times 3!4! \] then I divide that by the total number of possible arrangements of those seven letters, 7!

OpenStudy (anonymous):

Oh, wow satellite, you are right!

OpenStudy (anonymous):

when i did it i took the probability of xxEEExx as 1 1 3/5 2/4 1/3 1 1, but I didnt consider the chances the the first two letters WEREN'T Es

OpenStudy (anonymous):

ah right, and to think even simpler, (bonehead that i am) once we compute the probability that we get three in a row first, clearly we can ignore what happens next, because once you get three in a row the probability everything else gives a 1, that is once you have 3 in a row on the first slot, you can ignore the other combinations

OpenStudy (anonymous):

so with your thing satellite any 3-combo of the Es has a probability of 3/7 * 2/6 * 1/5 = 3/105

OpenStudy (anonymous):

there are 5 different possible 3-combos of E, so multiply that by 5 to get 3/21 = 1/7, but then we have to divide it by 7!/3! because that is the total number of possible outcomes, correct?

OpenStudy (anonymous):

It gives 1/5880

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i would say \[\frac{3}{105}=\frac{1}{35}\] and that is the probability you get them first. then you have 5 possible placements, they are disjoint events, so multiply by 5 and get \[\frac{1}{7}\]

OpenStudy (jamesj):

Yes, the answer is 1/7

OpenStudy (anonymous):

don't you still have to divide by the total number of outcomes?

OpenStudy (anonymous):

the difference between our methods is i am not concerned with how many possible ways i can pick the seven letters. i am just computing one probability. no, i accounted for that. if you want to divide by the number of possible outcomes, then use jamesj's method, which might make more sense

OpenStudy (anonymous):

if you want a 7! in the denominator your numerator has to include all the possible ways to get 3 E in a row

OpenStudy (jamesj):

Right. Sat has calculated the probability directly of each outcome EEExxxx, etc. and then added them. I counted the number of such outcomes, EEExxxx etc., then divided it by the total number of possible outcomes.

OpenStudy (anonymous):

I still don't understand.

OpenStudy (jamesj):

Let me ask you: how many ways can the letters be chosen to be in order EEELMNT

OpenStudy (anonymous):

OH I see, James, you did 5 x 3! divided by (7!/3!) didnt you?

OpenStudy (anonymous):

6 ways

OpenStudy (jamesj):

No. I have \[ \frac{5 3! 4!}{7!} \] === Yes 6 = 3!. Hence of all possible outcomes EEExxxx there are 3! x 4! possbilities.

OpenStudy (jamesj):

**correction: 5 . 3! 4! / 7!

OpenStudy (anonymous):

what? where did you get 4!?

OpenStudy (jamesj):

Because there are 4! ways of arraigning the L, M, N and T

OpenStudy (jamesj):

*arranging

OpenStudy (anonymous):

okay, so what is 5.

OpenStudy (jamesj):

EEExxxx --> 3!4! possibilities xEEExxx --> 3!4! possibilities xxEEExx --> 3!4! possibilities xxxEEEx --> 3!4! possibilities xxxxEEE --> 3!4! possibilities hence 5 x 3!4! possibilities in total

OpenStudy (anonymous):

alright, then you divided by 7! why? I know there are seven letters, but if you want to find the total possible outcomes you have to divide by (number of repeated letters)! and so it would be 3!

OpenStudy (jamesj):

No, there are 7! possible arrangements precisely because we counted the EEELMNT as 6 possible arrangements, not 1.

OpenStudy (jamesj):

If we counted EEELMNT as 1, then I would agree the total is 7!/3!

OpenStudy (anonymous):

Oh, I see what you did, say we did count EEELMNT as 1, what would we do from there?

OpenStudy (jamesj):

Then EEExxxx would have 4! possibilities and in total EEExxxx ... xxxxEEE would have 5 . 4! possibilities.

OpenStudy (anonymous):

Then divide by...?

OpenStudy (jamesj):

...and the probability would be \[ \frac{ 5 \cdot 4!}{7!/3!} = \frac{ 5 \cdot 4! 3!}{7!} = \frac{1}{7} \]

OpenStudy (anonymous):

Okay. Thanks. Now will you tell me what what is wrong with my way please? I have to go eat so take your time

OpenStudy (jamesj):

I didn't see your method, I'm sorry. Post it again and when I'm back from answering other questions myself, I'll see if I can comment on it.

OpenStudy (anonymous):

Alright. Well, I knew there was 5 different outcomes of 3 Es, so i found the probability of each like EEExxxx, xEEExxx, xxEEExx, xxxEEEx, xxxxEEE by doing 3/7 * 2/6 * 1/5 3/6 * 2/5 * 1/4 3/5 * 2/4 * 1/3 3/4 * 2/3 * 1/2 1 * 1 * 1, then I added the 5 probabilities together and divided by (7!/3!). I know I found each set of Es probability wrong. For example on xxEEExx i put 3/5 * 2/4 * 1/3 but it should've been 4/7 * 3/6 * 3/5 * 2/4 * 1/3. But that would've gotten me even further away from 1/7, my way I got 1/588, after fixing that mistake I got 1/5880 by doing (1/7)/(7!/3!). If the answer is 1/7, then how come I do not divide there?

OpenStudy (jamesj):

I'm still confused what exactly is your method.

OpenStudy (anonymous):

which part?

OpenStudy (jamesj):

Ok, I think I see what's going on here now. Each of the probabilities you calculated is exactly that: a probability. Why are you then dividing them by some other number of outcomes. You have already normalized them; you have already put in a denominator in calculating probabilities. To do that again is unnecessary.

OpenStudy (anonymous):

Oh, so the exact probability of xEEExxx is 4/7 * 3/6 * 2/5 * 1/4 * 1 * 1 * 1 not all of that divided by 7!/3!, that makes sense now. Thanks for all of your help James. That means I missed that question yesterday. Oh well ,we still got first place as a team.